# The myth of declining competitive balance in the UEFA Champions League group stage

### v1

@@ 1 Introduction
Recent studies have found a robust decline in both the ex ante (Triguero-Ruiz and Avila-Cano,, 2023) and ex post (Ramchandani et al.,, 2023; Triguero-Ruiz and Avila-Cano,, 2023) competitive balance of the UEFA Champions League group stage over the last decades, which might explain its reform. Our paper aims to check the robustness of these findings because we think that there exist alternative but reasonable measures of competitive balance.
@@
Based on the above arguments, two indices of ex ante and four indices of ex post competitive balance will be constructed in order to implement a proper sensitivity analysis. In contrast to previous results, we find no evidence for any decline in the competitive balance of the UEFA Champions League group stage between the 2003/04 and 2023/24 seasons. Consequently, if UEFA has chosen the barely used incomplete round-robin design of the UEFA Champions League with its inherent risks (Devriesere et al.,, 2024) because of the worsening trend in competitive balance, the decision-makers might have been misled.
@@ 2 Methodology and data
Sections 2.1 and 2.2 motivate and define the proposed indices for ex ante and ex post competitive balance, respectively. Section 2.3 provides an example of their calculation, while Section 2.4 gives information about our data.
@@ Acknowledgements
The research was supported by the National Research, Development and Innovation Office under Grants FK 145838 and PD 146055.

### v2

@@ 1 Introduction
Recent studies have found a robust decline in both the ex ante (Triguero-Ruiz and Avila-Cano,, 2023) and ex post (Ramchandani et al.,, 2023; Triguero-Ruiz and Avila-Cano,, 2023) competitive balance of the UEFA Champions League group stage over the last decades, which might explain its reform. Our paper aims to check the robustness of these findings: we think that there exist alternative but reasonable measures of competitive balance.
@@
Based on the above arguments, two indices of ex ante and four indices of ex post competitive balance will be constructed in order to implement a proper sensitivity analysis. In contrast to previous results, we find no evidence of any decline in the competitive balance of the UEFA Champions League group stage between the 2003/04 and 2023/24 seasons. Consequently, if UEFA has chosen the barely used incomplete round-robin design of the UEFA Champions League with its inherent risks (Devriesere et al.,, 2024) because of the worsening trend in competitive balance, the decision-makers might have been misled.
2 Related literature
The economic analysis of team sports is often centred around the issue of competitive balance since the seminal articles of Rottenberg, (1956) and El-Hodiri and Quirk, (1971). Fort and Maxcy, (2003) distinguish two approaches. Analysis of competitive balance (ACB) focuses on what has happened to competitive balance over time or as a result of rule changes. The second line of literature on competitive balance studies its effect on the stakeholders, especially fans, and usually involves testing the longstanding uncertainty of outcome hypothesis (UOH).
Measuring competitive balance in sports has evolved, with various metrics developed to quantify this crucial aspect of sporting contests. The survey of Pawlowski and Nalbantis, (2019) classifies the known measures according to the dimension of time. Short-term competitive balance refers to the uncertainty of individual games, often measured by using winning probabilities. The mid-term dimension focuses on seasonal uncertainty, while the long-term dimension concentrates on inter-seasonal uncertainty, the trend of competitive balance across several seasons. Our paper deals only with long-term competitive balance.
One commonly used measure of long-term competitive balance is the dispersion of winning percentage within sports leagues, also known as the Noll–Scully ratio (Noll,, 1988; Scully,, 1989). The Competitive Balance Ratio (CBR) proposed by Humphreys, (2002) is an alternative measure that captures season-to-season changes in relative standings.
The Noll–Scully ratio is biased when leagues with different season lengths are compared. To address this issue, Doria and Nalebuff, (2021) developed new metrics that are invariant to season length. The new measure has led to the reevaluation of competitive balance rankings among some major sports leagues, with surprising results such as the NFL moving from the most balanced to the least balanced league. Our study offers a similar re-evaluation of a “common belief” on competitive balance.
The other popular concept to measure long-term competitive balance is the Herfindahl–Hirschman index from the literature of industrial organization (Herfindahl,, 1950; Hirschman,, 1945, 1964), firstly used in sport by Depken, (1999). The normalized version of HHI is even more widespread (Owen et al.,, 2007). Triguero Ruiz and Avila-Cano, (2019) construct a novel index based on the concept of distance, with a range in the unit interval, which is the square root of the normalized HHI. It solves the problem of limited cardinality of most indices and makes the interpretation of the differences between the levels of competitive balance possible. Avila-Cano et al., (2021) determine its maximum concentration as the function of the number of teams and the scoring system. Analogously, Triguero-Ruiz et al., (2023) compute the minimal HHI in sports leagues without ties.
Naturally, some alternative measures of long-term competitive balance have also been suggested. Manasis et al., (2013) introduce a new measure to capture the degree of competition for winning the important prizes awarded in the European football leagues such as the championship, the qualification for lucrative international club competitions, and avoiding relegation. Csató and Petróczy, (2024) demonstrate the use of bibliometric indices for quantifying competitive balance in the UEFA Champions League knockout stage. A recent empirical analysis of nontransitive patterns in professional football (van Ours,, 2024) can also provide a novel approach to measuring competitive balance.
@@ 3 Methodology and data
Sections 3.1 and 3.2 motivate and define the proposed indices for ex ante and ex post competitive balance, respectively. Section 3.3 provides an example of their calculation, while Section 3.4 gives information about our data.
@@ Acknowledgements
The research was supported by the National Research, Development and Innovation Office under Grants FK 145838 and PD 146055, and by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.